# Turbocharging Gaussian Process Inference with Approximate Sketch-and-Project

### v1

@@ Abstract
Gaussian processes (GPs) play an essential role in biostatistics, scientific machine learning, and Bayesian optimization for their ability to provide probabilistic predictions and model uncertainty. However, GP inference struggles to scale to large datasets (which are common in modern applications), since it requires the solution of a linear system whose size scales quadratically with the number of samples in the dataset. We propose an approximate, distributed, accelerated sketch-and-project algorithm (ADASAP) for solving these linear systems, which improves scalability. We use the theory of determinantal point processes to show that the posterior mean induced by sketch-and-project rapidly converges to the true posterior mean. In particular, this yields the first efficient, condition number-free algorithm for estimating the posterior mean along the top spectral basis functions, showing that our approach is principled for GP inference. ADASAP outperforms state-of-the-art solvers based on conjugate gradient and coordinate descent across several benchmark datasets and a large-scale Bayesian optimization task. Moreover, ADASAP scales to a dataset with $>3\cdot 10^{8}$ samples, a feat which has not been accomplished in the literature.
@@ 1 Introduction
To address this gap in the literature, we introduce the Approximate Distributed Accelerated Sketch-and-Project (ADASAP) algorithm. ADASAP is rooted in the sketch-and-project framework of Gower and Richtárik (2015), and obtains the robustness to ill-conditioning of PCG and the scalability of SDD, while having reliable, default hyperparameters that work well in practice. For example, ADASAP dramatically outperforms tuned SDD and PCG on a dataset with $n>10^{6}$ samples (Fig. 1).
@@
    Using the theory of determinantal point processes, we show that sketch-and-project-style methods converge faster than stochastic first-order methods like SDD in the presence of ill-conditioning. In particular, we give a first-of-its-kind condition number-free time complexity bound for GP posterior mean estimation.
@@
Figure 1: ADASAP attains lower root mean square error (RMSE) and mean negative log likelihood (NLL) than start-of-the-art methods SDD (Lin et al., 2024) and PCG on the houseelec dataset. The solid lines indicate the mean performance of each method, while the shaded regions indicate the range between worst and best performance of each method over five random splits of the data.
@@ 2.1 Pathwise conditioning
Equation (4) enables scalable sampling from the posterior as $f$ can be approximated using random features (Rahimi and Recht, 2007; Rudi and Rosasco, 2017) for $k$, which costs $\mathcal{O}(q|X_{*}|)$, where $q$ is the number of random features. This yields a significant reduction in cost from the standard posterior sampling scheme in (3): by leveraging pathwise conditioning, posterior sampling can be performed by generating $f$ via random features (Wilson et al., 2020, 2021) and solving $(K+\lambda I)^{-1}(y-f(X_{n})-\zeta)$ using an iterative method like PCG or SDD. Our experiments (Section 5) use pathwise conditioning to sample from the posterior.
@@ 2.2 Drawbacks of state-of-the-art exact inference methods
The current state-of-the-art for large-scale posterior sampling is SDD (Lin et al., 2024), a stochastic first-order algorithm combining coordinate descent (Richtárik and Takáč, 2014; Qu and Richtárik, 2016) with momentum and geometric averaging. Stochastic methods like SDD can better handle single (or lower) precision than PCG (Wu et al., 2024; Rathore et al., 2025), which is essential to maximizing GPU performance and reducing storage costs (Abdelfattah et al., 2021). While SDD overcomes the scalability challenges associated with PCG, it faces challenges due to the ill-conditioning of $K+\lambda I$, which determines its worst-case convergence rate (Nemirovski and Yudin, 1983). Therefore, reliable large-scale inference requires a method that has cheaper per-iteration costs than PCG and can handle the ill-conditioning of $K+\lambda I$. However, to the best of our knowledge, no existing method effectively tackles both issues: SGD and SDD offer speed and scalability but suffer from ill-conditioning, while PCG offers robustness to ill-conditioning but struggles to scale beyond $n>10^{6}$. In Section 3 we propose SAP, which is robust to ill-conditioning; in Section 4, we develop ADASAP, an extension of SAP that scales to large datasets.
@@ 3 Sketch-and-project for GP Inference
Section 2 shows the main challenge of large-scale GP inference is the need to solve a large linear system with the ill-conditioned matrix $K+\lambda I$, which cannot be done satisfactorily with PCG or SDD. We introduce the SAP algorithm, which effectively addresses the conditioning challenges of GP inference. However, SAP alone does not address the scalability challenges of GP inference, motivating the development of ADASAP in Section 4.
@@ 3.1 Sketch-and-project
distribution of row indices $\mathcal{D}$ over $[n]^{b}$, number of iterations $T_{\max}$, initialization $W_{0}\in\mathbb{R}^{n\times n_{\textrm{rhs}}}$, averaging boolean tail_average
@@
     $D_{t}\leftarrow I_{\mathcal{B}}^{T}(K_{\mathcal{B}\mathcal{B}}+\lambda I)^{-1}(K_{\mathcal{B}n}W_{t}+\lambda I_{\mathcal{B}}W_{t}-Y_{\mathcal{B}})$triangleright Search direction; costs $\mathcal{O}(bnn_{\textrm{rhs}}+b^{3})$
     $W_{t+1}\leftarrow W_{t}-D_{t}$triangleright Update parameters; costs $\mathcal{O}(bn_{\textrm{rhs}})$
@@
return $\bar{W}_{T}$ if tail_average else $W_{T}$triangleright Approximate solution to $K_{\lambda}W=Y$
We formally present SAP in Algorithm 1. SAP resembles randomized coordinate descent, except that it employs subspace preconditioning. That is, it preconditions the subspace gradient $K_{\mathcal{B}n}W_{t}+\lambda I_{\mathcal{B}}W_{t}-Y_{\mathcal{B}}$ by the inverse of the subspace Hessian $K_{\mathcal{B}\mathcal{B}}+\lambda I$. Thus, unlike SDD, SAP includes second-order information in its search direction. In addition, we have also augmented vanilla SAP from Gower and Richtárik (2015) with tail averaging (Jain et al., 2018; Epperly et al., 2024a), which helps address the inherent noise in the iterates, and plays a crucial role in our analysis.
SAP (assuming $n_{\textrm{rhs}}=1$ for simplicity) only costs $\mathcal{O}(bn+b^{3})$ per-iteration, which is significantly smaller than the $\mathcal{O}(n^{2})$ iteration cost of PCG, while being comparable with SDD, which costs $\mathcal{O}(bn)$ per-iteration. Furthermore, PCG typically requires memory linear in $n$ to store the preconditioner, while SAP only incurs and additional $\mathcal{O}(b^{2})$ storage cost. Thus, SAP strikes a balance between PCG and methods like SDD. It incorporates second-order information, like PCG, but does so at a modest increase in cost relative to SDD and SGD. By leveraging the theory of determinantal point processes (Kulesza and Taskar, 2012), we will show tail-averaged SAP improves dependence upon the condition number, which establishes it as the solution to the conditioning challenges discussed in Section 2.
3.2 Fast convergence along $-\ell$ subspace for GP inference
Prior works (Dicker et al., 2017; Lin et al., 2020, 2023) have shown that the components of the solution most relevant for learning lie along the dominant eigenvectors of $K$, which correspond to the dominant spectral basis functions of the reproducing kernel Hilbert space (RKHS) $\mathcal{H}$ induced by kernel $k$ with respect to $X$. Lin et al. (2023) shows that SGD enjoys an improved early convergence along these directions, but attains a slow sublinear rate asymptotically. Motivated by these observations, we study the convergence rate of tail-averaged SAP along the dominant spectral basis functions, showing that it obtains a two-phase convergence rate: an initial condition number-free, sublinear rate, and an asymptotic linear rate controlled by the blocksize $b$.
@@ 3.2 Fast convergence along $-\ell$ subspace for GP inference
We focus on posterior mean estimation, which corresponds to solving the linear system $K_{\lambda}W=y$ using tail-averaged SAP. Recall that, given a sequence of weight vectors $W_{1},...,W_{t}$ produced by sketch-and-project (which are vectors, not matrices, as $n_{\textrm{rhs}}=1$), tail averaging obtains a weight estimate by averaging the second half. We use these weights to define an estimate of the GP posterior mean $m_{n}$, which is given by $\hat{m}_{t}=k(\cdot,X)\bar{W}_{t}$.
Theorem 3.2 is our main theoretical result. It provides a convergence guarantee for the SAP posterior mean estimate $\hat{m}_{t}$ along the $-\ell$ spectral basis functions, where $\ell$ is a parameter that can be chosen freely. Theorem 3.2 shows that while the estimate converges to the true posterior mean over the entire RKHS $\mathcal{H}$, its initial convergence rate along the dominant basis functions is even faster. Our theoretical results rely on the smoothed condition number, which we define below.
@@
The smoothed condition number controls both the SGD-style sublinear convergence rate and the PCG-style linear convergence rate in Theorem 3.2. Note that, for notational convenience, we use $2b$ to denote the SAP blocksize in the below statement.
@@
Suppose we have a kernel matrix $K\in\mathbb{R}^{n\times n}$, observations $y\in\mathbb{R}^{n}$ from a GP prior with likelihood variance $\lambda\geq 0$, and let $\phi(\cdot,\cdot)$ denote the smoothed condition number of $K+\lambda I$. Given any $\ell\in[n]$, let $\mathrm{proj}_{\ell}$ denote orthogonal projection onto the span of $u^{(1)},...,u^{(\ell)}$. For any $21\leq b\leq n/$ 2, in $\tilde{\mathcal{O}}(nb^{2})$ time we can construct a distribution over row index subsets $\mathcal{B}$ of size $2b$ such that SAP (Algorithm 1) initialized at zero after $t$ iterations satisfies
@@
For ill-conditioned kernels with small prior variance $\lambda$, the linear rate $1-1/(2\phi(b,n))$ close to 1 (as $b$ increases, the preconditioning effect becomes stronger, which improves the rate). On the other hand, the constant in the sublinear rate $\phi(b,\ell)$ is much smaller, since it takes advantage of the $-\ell$ subspace norm, which makes it effectively independent of the conditioning of $K$.
If the SAP blocksize is proportional to $\ell$ and $K_{\lambda}$ exhibits any polynomial spectral decay, then $\phi(b,\ell)=\mathcal{O}(1)$, making the sublinear rate condition number-free. Most popular kernels exhibit polynomial spectral decay (Ma and Belkin, 2017; Kanagawa et al., 2018). Furthermore, polynomial spectral decay is a standard assumption in the generalization analysis of kernel methods in the fixed and random design settings (Caponnetto and DeVito, 2007; Bach, 2013; Rudi et al., 2015). To the best of our knowledge, our result is the first of its kind, as previous comparable guarantees for stochastic methods (Lin et al., 2023, Proposition 1) still depend on the condition number. We illustrate this claim with the following corollary:
@@
The $\tilde{\mathcal{O}}((n^{2}+n\ell^{2})/\epsilon)$ runtime (in phase one) is entirely independent of the condition number of $K$ and the likelihood variance $\lambda$. In Appendix B, we show that when $\lambda$ is sufficiently small (the highly ill-conditioned case), then the phase one complexity of sketch-and-project can be further improved by increasing the block size $b$ to attain $\tilde{\mathcal{O}}(nb^{2}+(n^{2}+nb^{2})(\ell/b)^{\beta-1}/\epsilon)$ runtime. In particular, when $\ell\ll b\ll\sqrt{n}$, we obtain meaningful convergence in $o(n^{2})$ time, i.e., before reading all of the data. Thus, the $-\ell$ subspace convergence of SAP can actually benefit from ill-conditioned kernels as the leading spectral basis functions become more dominant.
@@ 4 ADASAP: Approximate, Distributed, Accelerated SAP
We introduce ADASAP (Algorithm 2), a scalable extension of SAP for GP inference. We first introduce the modifications made to SAP for scalability (Section 4.1), before presenting ADASAP in full (Section 4.2). For a detailed discussion of how ADASAP relates to prior work, please see Appendix A.
@@ 4.1 The key ingredients: approximation, distribution, and acceleration
Approximate subspace preconditioning
SAP enjoys significant improvements over PCG, as it only requires $\mathcal{O}(bn+b^{3})$ computation per-iteration. Unfortunately, SAP is limited in how large a blocksize $b$ it may use, as factoring $K_{\mathcal{B}\mathcal{B}}+\lambda I$ to perform subspace preconditioning costs $\mathcal{O}(b^{3})$. This is problematic, as Theorem 3.2 shows the convergence rate of SAP improves as $b$ increases. To address this challenge, ADASAP draws inspiration from previous works (Erdogdu and Montanari, 2015; Frangella et al., 2024b; Rathore et al., 2025), which replace exact linear system solves in iterative algorithms with inexact solves based on low-rank approximations. ADASAP replaces $K_{\mathcal{B}\mathcal{B}}$ in SAP with a rank-$r$ randomized Nyström approximation $\hat{K}_{\mathcal{B}\mathcal{B}}$ (Williams and Seeger, 2000; Gittens and Mahoney, 2016; Tropp et al., 2017). This strategy is natural, as kernels exhibit approximate low-rank structure (Bach, 2013; Rudi et al., 2015; Belkin, 2018). Indeed, Rathore et al. (2025) develops an approximate SAP solver for kernel ridge regression with one right-hand side, based on approximating $K_{\mathcal{B}\mathcal{B}}$ by $\hat{K}_{\mathcal{B}\mathcal{B}}$, and observes strong empirical performance. ADASAP computes $\hat{K}_{\mathcal{B}\mathcal{B}}$ following the numerically stable procedure from Tropp et al. (2017), which is presented in Appendix C. The key benefit is $(\hat{K}_{\mathcal{B}\mathcal{B}}+\rho I)^{-1}$ can be applied to vectors in $\mathcal{O}(br)$ time (Appendix C) compared to $\mathcal{O}(b^{3})$ for the exact SAP update. This reduction allows ADASAP to use larger blocksizes: on the taxi dataset ($n=3.31\cdot 10^{8}$), ADASAP uses blocksize $b=1.65\cdot 10^{5}$.
Distributed matrix-matrix products
SAP with approximate subspace preconditioning allows large blocksize $b$, but two bottlenecks remain: computing $K_{\mathcal{B}n}W_{t}$ and constructing the sketch $K_{\mathcal{B}\mathcal{B}}\Omega$. The former costs $\mathcal{O}(bn)$, while the latter costs $\mathcal{O}(b^{2}r)$. As matrix-matrix multiplication is embarrassingly parallel, distributed multi-GPU acceleration can address these bottlenecks. By partitioning $K_{\mathcal{B}n}$ across $\mathrm{N_{workers}}$ GPUs, ColDistMatMat significantly reduces the time to compute the product $K_{\mathcal{B}n}\Omega$. Similarly, RowDistMatMat for $K_{\mathcal{B}\mathcal{B}}\Omega$ reduces the time to compute the sketch. On our largest dataset, taxi, using 4 GPUs achieves a 3.4×3.4$\times$3.4 × speedup (Fig. 4), which corresponds to near-perfect parallelism.
Nesterov acceleration

### v2

@@ Abstract
Gaussian processes (GPs) play an essential role in biostatistics, scientific machine learning, and Bayesian optimization for their ability to provide probabilistic predictions and model uncertainty. However, GP inference struggles to scale to large datasets (which are common in modern applications), since it requires the solution of a linear system whose size scales quadratically with the number of samples in the dataset. We propose an approximate, distributed, accelerated sketch-and-project algorithm (ADASAP) for solving these linear systems, which improves scalability. We use the theory of determinantal point processes to show that the posterior mean induced by sketch-and-project rapidly converges to the true posterior mean. In particular, this yields the first efficient, condition number-free algorithm for estimating the posterior mean along the top spectral basis functions, showing that our approach is principled for GP inference. ADASAP outperforms state-of-the-art solvers based on conjugate gradient and coordinate descent across several benchmark datasets and a large-scale Bayesian optimization task. Moreover, ADASAP scales to a dataset with $3\cdot 10^{8}$ samples, a feat which has not been accomplished in the literature.
@@ 1 Introduction
To address this gap in the literature, we introduce the Approximate Distributed Accelerated Sketch-and-Project (ADASAP) algorithm. ADASAP is rooted in the sketch-and-project framework of Gower and Richtárik (2015), and obtains the robustness to ill-conditioning of PCG and the scalability of SDD, while having reliable, default hyperparameters that work well in practice. For example, ADASAP dramatically outperforms tuned SDD and PCG on a dataset with $n>10^{6}$ samples (Fig.˜1).
@@
    Using the theory of determinantal point processes, we show that sketch-and-project-style methods converge faster than stochastic first-order methods like SDD in the presence of ill-conditioning. In particular, we give a first-of-its-kind condition number-free time complexity bound for estimating the GP posterior mean to moderate accuracy, explaining the excellent test error performance of ADASAP.
@@
Figure 1: ADASAP attains lower root mean square error (RMSE) and mean negative log likelihood (NLL) than start-of-the-art methods SDD (Lin et al., 2024) and PCG on the houseelec dataset. SDD-1 and SDD-10 correspond to two particular stepsize selections for the SDD method. The solid lines indicate the mean performance of each method, while the shaded regions indicate the range between worst and best performance of each method over five random splits of the data.
@@ 2.1 Pathwise conditioning
Equation (4) enables scalable sampling from the posterior as $f$ can be approximated using random features (Rahimi and Recht, 2007; Rudi and Rosasco, 2017) for $k$, which costs $\mathcal{O}(q|X_{*}|)$, where $q$ is the number of random features. This yields a significant reduction in cost from the standard posterior sampling scheme in (3): by leveraging pathwise conditioning, posterior sampling can be performed by generating $f$ via random features (Wilson et al., 2020, 2021) and solving $(K+\lambda I)^{-1}(y-f(X_{n})-\zeta)$ using an iterative method like PCG or SDD. Our experiments (Section˜5) use pathwise conditioning to sample from the posterior.
@@ 2.2 Drawbacks of state-of-the-art exact inference methods
The current state-of-the-art for large-scale posterior sampling is SDD (Lin et al., 2024), a stochastic first-order algorithm combining coordinate descent (Richtárik and Takáč, 2014; Qu and Richtárik, 2016) with momentum and geometric averaging. Stochastic methods like SDD can better handle single (or lower) precision than PCG (Wu et al., 2024; Rathore et al., 2025), which is essential to maximizing GPU performance and reducing storage costs (Abdelfattah et al., 2021). While SDD overcomes the scalability challenges associated with PCG, it faces challenges due to the ill-conditioning of $K+\lambda I$, which determines its worst-case convergence rate (Nemirovski and Yudin, 1983). Therefore, reliable large-scale inference requires a method that has cheaper per-iteration costs than PCG and can handle the ill-conditioning of $K+\lambda I$. However, to the best of our knowledge, no existing method effectively tackles both issues: SGD and SDD offer speed and scalability but suffer from ill-conditioning, while PCG offers robustness to ill-conditioning but struggles to scale beyond $n>10^{6}$. In Section˜3 we propose SAP, which is robust to ill-conditioning; in Section˜4, we develop ADASAP, an extension of SAP that scales to large datasets.
@@ 3 Sketch-and-project for GP Inference
Section˜2 shows the main challenge of large-scale GP inference is the need to solve a large linear system with the ill-conditioned matrix $K+\lambda I$, which cannot be done satisfactorily with PCG or SDD. We introduce the SAP algorithm, which effectively addresses the conditioning challenges of GP inference. However, SAP alone does not address the scalability challenges of GP inference, motivating the development of ADASAP in Section˜4.
@@ 3.1 Sketch-and-project
blocksize $b$, distribution of row indices $\mathcal{D}$ over $n]^{b}$, number of iterations $T_{\max}$, initialization $W_{0}\in\mathbb{R}^{n\times n_{\textrm{rhs}}}$, averaging boolean tail_average
@@
  $D_{t}\leftarrow I_{\mathcal{B}}^{T}(K_{\mathcal{B}\mathcal{B}}+\lambda I)^{-1}(K_{\mathcal{B}n}W_{t}+\lambda I_{\mathcal{B}}W_{t}-Y_{\mathcal{B}})$ $\triangleright$ Search direction; costs $\mathcal{O}(bnn_{\textrm{rhs}}+b^{3})$
  $W_{t+1}\leftarrow W_{t}-D_{t}$ $\triangleright$ Update parameters; costs $\mathcal{O}(bn_{\textrm{rhs}})$
@@
return $\bar{W}_{T}$ if tail_average else $W_{T}$ $\triangleright$ Approximate solution to $K_{\lambda}W=Y$
We formally present SAP in Algorithm˜1. SAP resembles randomized block coordinate descent, except that it employs subspace preconditioning. That is, it preconditions the subspace gradient $K_{\mathcal{B}n}W_{t}+\lambda I_{\mathcal{B}}W_{t}-Y_{\mathcal{B}}$ by the inverse of the subspace Hessian $K_{\mathcal{B}\mathcal{B}}+\lambda I$. Thus, unlike SDD, SAP includes second-order information in its search direction. In addition, we have also augmented vanilla SAP from Gower and Richtárik (2015) with tail averaging (Jain et al., 2018; Epperly et al., 2024a), which helps address the inherent noise in the iterates, and plays a crucial role in our analysis.
SAP (assuming the number of right-hand sides $n_{\textrm{rhs}}=1$ for simplicity) only costs $\mathcal{O}(bn+b^{3})$ per-iteration, which is significantly smaller than the $\mathcal{O}(n^{2})$ iteration cost of PCG, while being comparable with SDD, which costs $\mathcal{O}(bn)$ per-iteration. Furthermore, PCG typically requires memory linear in $n$ to store the preconditioner, while SAP only incurs and additional $\mathcal{O}(b^{2})$ storage cost. Thus, SAP strikes a balance between PCG and methods like SDD. It incorporates second-order information, like PCG, but does so at a modest increase in cost relative to SDD and SGD. By leveraging the theory of determinantal point processes (Kulesza and Taskar, 2012), we show that tail-averaged SAP improves dependence upon the condition number in the early phase of the convergence, which establishes it as a potential solution to the conditioning challenges discussed in Section˜2, particularly in the presence noisy data and generalization error.
3.2 Fast convergence along top-$\ell$ subspace for GP inference
Prior works (Dicker et al., 2017; Lin et al., 2020, 2023) have shown that the components of the solution most relevant for learning lie along the dominant eigenvectors of $K$, which correspond to the dominant spectral basis functions of the reproducing kernel Hilbert space (RKHS) $\mathcal{H}$ induced by kernel $k$ with respect to $X$. Lin et al. (2023) shows that SGD enjoys an improved early convergence along these directions, but attains a slow sublinear rate asymptotically. Motivated by these observations, we study the convergence rate of tail-averaged SAP along the dominant spectral basis functions, showing that it obtains a two-phase convergence rate: an initial condition number-free sublinear rate, followed by an asymptotic linear rate where the condition number dependence is mitigated by the blocksize $b$. The proofs of the results in this section appear in Section˜B.3.
@@ 3.2 Fast convergence along top-$\ell$ subspace for GP inference
We focus on posterior mean estimation, which corresponds to solving the linear system $K_{\lambda}W=y$ using tail-averaged SAP. Recall that, given a sequence of weight vectors $W_{1},...,W_{t}$ produced by sketch-and-project (which are vectors, not matrices, as $n_{\textrm{rhs}}=1$), tail averaging obtains a weight estimate by averaging the second half of this sequence. We use these weights to define an estimate of the GP posterior mean $m_{n}$, which is given by $\hat{m}_{t}=k(\cdot,X)\bar{W}_{t}$.
Theorem˜3.2 is our main theoretical result. It provides a convergence guarantee for the SAP posterior mean estimate $\hat{m}_{t}$ along the top-$\ell$ spectral basis functions, where $\ell$ is a parameter that can be chosen freely. Theorem˜3.2 shows that while the estimate converges to the true posterior mean over the entire RKHS $\mathcal{H}$, its initial convergence rate along the dominant basis functions is even faster. Our theoretical results rely on the smoothed condition number, which we define below.
@@
The smoothed condition number controls both the SGD-style sublinear convergence rate and the PCG-style linear convergence rate in Theorem˜3.2. Note that, for notational convenience, we use $2b$ to denote the SAP blocksize in the below statement.
@@
Suppose we have a kernel matrix $K\in\mathbb{R}^{n\times n}$, observations $y\in\mathbb{R}^{n}$ from a GP prior with likelihood variance $\lambda\geq 0$, and let $\phi(\cdot,\cdot)$ denote the smoothed condition number of $K+\lambda I$. Given any $\ell\in[n]$, let $\mathrm{proj}_{\ell}$ denote orthogonal projection onto the span of $u^{(1)},...,u^{(\ell)}$. For any $1\leq b\leq n/2$, in $\tilde{\mathcal{O}}(nb^{2})$ time we can construct a distribution over row index subsets $\mathcal{B}$ of size $2b$ such that SAP (Algorithm 1) initialized at zero after $t$ iterations satisfies
@@
For ill-conditioned kernels with small prior variance $\lambda$, the linear rate $1-1/(2\phi(b,n))$ is close to 1 (as $b$ increases, the preconditioning effect becomes stronger, which improves the rate). On the other hand, the constant in the sublinear rate $\phi(b,\ell)$ is much smaller, since it takes advantage of the top-$\ell$ subspace norm, which makes it effectively independent of the conditioning of $K$.
If the SAP blocksize is proportional to $\ell$ and $K+\lambda I$ exhibits any polynomial spectral decay, then $\phi(b,\ell)=\mathcal{O}(1)$, making the sublinear rate condition number-free. Most popular kernels exhibit polynomial spectral decay (Ma and Belkin, 2017; Kanagawa et al., 2018). Furthermore, polynomial spectral decay is a standard assumption in the generalization analysis of kernel methods in the fixed and random design settings (Caponnetto and DeVito, 2007; Bach, 2013; Rudi et al., 2015). To the best of our knowledge, our result is the first of its kind, as previous comparable guarantees for stochastic methods (Lin et al., 2023, Proposition 1) still depend on the condition number. We note that our condition number-free rate is attained only when seeking moderately accurate estimates of the posterior mean (which is often the case due to the presence of noise in the data). When seeking highly accurate estimates, the asymptotic (condition number dependent) linear rate will eventually take over. We illustrate these claims with the following corollary:
@@
The $\tilde{\mathcal{O}}((n^{2}+n\ell^{2})/\epsilon)$ runtime (phase one, attained for moderate accuracy $\epsilon$) is entirely independent of the condition number of $K$ and the likelihood variance $\lambda$. In Appendix˜B, we show that when $\lambda$ is sufficiently small (the highly ill-conditioned case), then the phase one complexity of sketch-and-project can be further improved by increasing the block size $b$ to attain $\tilde{\mathcal{O}}(nb^{2}+(n^{2}+nb^{2})(\ell/b)^{\beta-1}/\epsilon)$ runtime. In particular, when $\ell\ll b\ll\sqrt{n}$, we obtain meaningful convergence in $o(n^{2})$ time, i.e., before reading all of the data. Thus, the top-$\ell$ subspace convergence of SAP can actually benefit from ill-conditioned kernels as the leading spectral basis functions become more dominant.
Although our theoretical results require tail averaging, we do not believe it is needed for good practical performance of sketch-and-project algorithms. Indeed, we run sketch-and-project with and without tail averaging in Section˜C.6, and we find that (i) tail averaging does not improve subspace convergence by a substantial amount and (ii) tail averaging performs worse than not using tail averaging at larger blocksizes.
@@ 4 ADASAP: Approximate, Distributed, Accelerated SAP
We introduce ADASAP (Algorithm˜2), a scalable extension of SAP for GP inference. We first introduce the modifications made to SAP for scalability (Section˜4.1), before presenting ADASAP in full (Section˜4.2). For a detailed discussion of how ADASAP relates to prior work, please see Appendix˜A.
@@ 4.1 The key ingredients: approximation, distribution, and acceleration
4.1.1 Approximate subspace preconditioning
SAP enjoys significant improvements over PCG, as it only requires $\mathcal{O}(bn+b^{3})$ computation per-iteration. Unfortunately, SAP is limited in how large a blocksize $b$ it may use, as factoring $K_{\mathcal{B}\mathcal{B}}+\lambda I$ to perform subspace preconditioning costs $\mathcal{O}(b^{3})$. This is problematic, as Theorem˜3.2 shows the convergence rate of SAP improves as $b$ increases.
To address this challenge, ADASAP draws inspiration from previous works (Erdogdu and Montanari, 2015; Frangella et al., 2024b; Rathore et al., 2025), which replace exact linear system solves in iterative algorithms with inexact solves based on low-rank approximations. ADASAP replaces $K_{\mathcal{B}\mathcal{B}}$ in SAP with a rank-$r$ randomized Nyström approximation $\hat{K}_{\mathcal{B}\mathcal{B}}$ (Williams and Seeger, 2000; Gittens and Mahoney, 2016; Tropp et al., 2017). This strategy is natural, as kernels exhibit approximate low-rank structure (Bach, 2013; Rudi et al., 2015; Belkin, 2018). Indeed, Rathore et al. (2025) develops an approximate SAP solver for kernel ridge regression with one right-hand side, based on approximating $K_{\mathcal{B}\mathcal{B}}$ by $\hat{K}_{\mathcal{B}\mathcal{B}}$, and observes strong empirical performance.
ADASAP computes $\hat{K}_{\mathcal{B}\mathcal{B}}$ following the numerically stable procedure from Tropp et al. (2017), which is presented in Appendix˜C. The key benefit is that $(\hat{K}_{\mathcal{B}\mathcal{B}}+\rho I)^{-1}$ can be applied to vectors in $\mathcal{O}(br)$ time (Appendix˜C), where $\rho>0$ is a damping parameter to ensure invertibility. This reduces the cost compared to the $\mathcal{O}(b^{3})$ exact SAP update and allows ADASAP to use larger blocksizes: on the taxi dataset ($n=3.31\cdot 10^{8}$), ADASAP uses blocksize $b=1.65\cdot 10^{5}$.
4.1.2 Distributed matrix-matrix products
SAP with approximate subspace preconditioning allows large blocksize $b$, but two bottlenecks remain: computing $K_{\mathcal{B}n}W_{t}$ and constructing the sketch $K_{\mathcal{B}\mathcal{B}}\Omega$. The former costs $\mathcal{O}(bn)$, while the latter costs $\mathcal{O}(b^{2}r)$. As matrix-matrix multiplication is embarrassingly parallel, distributed multi-GPU acceleration can address these bottlenecks. By partitioning $K_{\mathcal{B}n}$ across $\mathrm{N_{workers}}$ GPUs, ColDistMatMat significantly reduces the time to compute the product $K_{\mathcal{B}n}\Omega$. Similarly, RowDistMatMat for $K_{\mathcal{B}\mathcal{B}}\Omega$ reduces the time to compute the sketch. On our largest dataset, taxi, using 4 GPUs achieves a $3.4\times$ speedup (Fig.˜4), which corresponds to near-perfect parallelism.
4.1.3 Nesterov acceleration